OpenStudy (anonymous):

there were 40,000 people in jhonson city.then the tv factory cloesed down. 8,000 people left town. what percent of the people left town?

OpenStudy (anonymous):

# left / # people 8000/40000 = 1/5 = 20%

OpenStudy (anonymous):

show me in long terms

OpenStudy (anonymous):

long terms?

OpenStudy (anonymous):

meverett showed all the work